# Learning Collective Variables with Synthetic Data Augmentation through Physics-inspired Geodesic Interpolation

### v3

@@ 3.4 Current limitations of ML CVs
As we currently do not have a satisfying answer to this problem, future work will have to find a way to increase the robustness of ML-based CVs in order to ensure the trustworthiness of the results obtained with them. A possible avenue might be training multi-task CVs as suggested by Bonati et al. 19, e.g., one where an auto-encoder framework is trained with a reconstruction loss combined with a TDA-type loss on the latent space. The reconstruction loss might introduce necessary robustness and is perfectly usable with synthetic, unlabeled data (lackign clear state labels) as provided by our geodesic interpolation method. Another approach might involve the inclusion of true transition data, as done in TPI-Deep-TDA 24, to guide the model in the correct direction. However, this would not align with the fast, one-shot approach of this work.
@@ 4 Conclusion
We developed a scheme to train collective variables (CVs) from metastable state configurations using geodesic interpolation to generate a synthetic transition state ensemble (TSE). We found that the generated TSE closely aligns with the reference transition configurations and provides additional information of the transition progress via the interpolation parameter $t$. By sampling the parameter from a predefined distribution, we could control the distribution of the generated TSE over the transition progress. Furthermore, we devised a method to reverse calculate the interpolation parameter, potentially enabling the incorporation of transition data obtained, e.g., by flooding simulations. By leveraging the interpolation parameter label as a regression target for CVs, we demonstrated notable improvements in the enhanced sampling simulation results when compared to the CVs trained solely on metastable state conformations. These results highlight the importance of incorporating transition state information into CV model training. This not only enhances the ability to accurately distinguish between the metastable states, but also enables us to effectively capture the underlying dynamics of the system, as evidenced by the change in a number of hydrogen bonds along the CV axis. We note that the behavior of ML CVs for out-of-distribution samples in enhanced sampling simulations could not be perfectly controlled with the use of pseudo-TSE data.
@@
The code to reproduce this work is available on GitHub: https://github.com/learningmatter-mit/geodesic-interpolation-cv 45. The PLUMED 46 input files are also available on PLUMED-NEST (www.plumed-nest.org), the public repository of the PLUMED consortium 47, as plumID:24.014.
@@
The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the research results reported in this article. The authors thank D. E. Shaw Research for sharing the reference trajectory data of chignolin (CLN025) in Lindorff-Larsen et al. 32. We also want to thank Martin Šípka, Luigi Bonati, and Gianmarco Lazzeri for helpful discussions. We acknowledge the anonymous reviewers for their valuable insights, including their suggestion to references 27, 19, 26, 21, 22. S.Y. is supported by Takeda Fellowship and Ilju Overseas PhD Fellowship. J.N. is supported by the Ronald A. Kurtz Fellowship. J.C.B.D. is thankful for the support of the Leopoldina Fellowship Program, German National Academy of Sciences Leopoldina, grant number LPDS 2021-08.
@@
The SI contains: i) details of the molecular dynamics and enhanced sampling simulations; ii) specifics of the data augmentation using geodesic interpolation and CV training; iii) structure-based analysis using hydrogen bonds present in the folded state; iv) additional simulation analysis, including the time evolution of CV and Cα TICA projections; v) analysis of computational costs.

### v4

@@ 3.4 Current limitations of ML CVs
As we currently do not have a satisfying answer to this problem, future work will have to find a way to increase the robustness of ML-based CVs in order to ensure the trustworthiness of the results obtained with them. A possible avenue might be training multitask CVs as suggested by Bonati et al. 19, e.g., one where an autoencoder framework is trained with a reconstruction loss combined with a TDA-type loss on the latent space. The reconstruction loss might introduce necessary robustness and is perfectly usable with synthetic, unlabeled data (lacking clear state labels) as provided by our geodesic interpolation method. Another approach might involve the inclusion of true transition data, as done in TPI-Deep-TDA 24, to guide the model in the correct direction. However, this would not align with the fast, one-shot approach of this work.
@@ 4 Conclusion
We developed a scheme to train collective variables (CVs) from metastable state configurations using geodesic interpolation to generate a synthetic transition state ensemble (TSE). We found that the generated TSE closely aligns with the reference transition configurations and provides additional information on the transition progress via the interpolation parameter $t$. By sampling the parameter from a predefined distribution, we could control the distribution of the generated TSE over the transition progress. Furthermore, we devised a method to reverse calculate the interpolation parameter, potentially enabling the incorporation of transition data obtained, e.g., by flooding simulations. By leveraging the interpolation parameter label as a regression target for CVs, we demonstrated notable improvements in the enhanced sampling simulation results when compared to the CVs trained solely on metastable state conformations. These results highlight the importance of incorporating transition state information into CV model training. This not only enhances the ability to accurately distinguish between the metastable states, but also enables us to effectively capture the underlying dynamics of the system, as evidenced by the change in a number of hydrogen bonds along the CV axis. We note that the behavior of ML CVs for out-of-distribution samples in enhanced sampling simulations could not be perfectly controlled with the use of pseudo-TSE data.
@@
The code to reproduce this work is available on GitHub: https://github.com/learningmatter-mit/geodesic-interpolation-cv. The PLUMED 45 input files are also available on PLUMED-NEST (www.plumed-nest.org), the public repository of the PLUMED consortium 46, as plumID:24.014.
@@
The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the research results reported in this article. The authors thank D. E. Shaw Research for sharing the reference trajectory data of chignolin (CLN025) in Lindorff-Larsen et al. 32. We also want to thank Martin Šípka, Luigi Bonati, and Gianmarco Lazzeri for helpful discussions. We acknowledge the anonymous reviewers for their valuable insights, including their suggestions to refs 27, 19, 26, 21, 22. S.Y. is supported by Takeda Fellowship and Ilju Overseas Ph.D. Fellowship. J.N. is supported by the Ronald A. Kurtz Fellowship. J.C.B.D. is thankful for the support of the Leopoldina Fellowship Program, German National Academy of Sciences Leopoldina, grant number LPDS 2021-08.
@@
Details of the molecular dynamics and enhanced sampling simulations, specifics of the data augmentation using geodesic interpolation and CV training, structure-based analysis using hydrogen bonds present in the folded state, additional simulation analysis, including the time evolution of CV and Cα TICA projections, and analysis of computational costs